find an equation of the line having the given slope and containing the given point. m=5/6, (5,-9).
The equation of the line is y=??

Answer by ewatrrr(24785)


Hi
line having the given slope and containing the given point. m=5/6, (5,-9).


y = (5/6)x + b |using ordered pair Pt(5,-9) to solve for b
-9 = 25/6 + b
-54/6 - 25/6 = b
-79/6 = b
y = (5/6)x -79/6